find the surface area of the cylinder in terms of 3.14 the radius is 9cm and the height is 20 cm
Answer by checkley77(12844)

CIRCUMFERENCE=2PIR=2*9PI=18PI
20*18PI=360PI IS THE TOTAL SURFACE AREA.